# LLM2Vec-Gen: Generative Embeddings from Large Language Models

### v1

@@ Abstract
LLM-based text embedders typically encode the semantic content of their input. However, embedding tasks require mapping diverse inputs to similar outputs. Typically, this input-output is addressed by training embedding models with paired data using contrastive learning. In this work, we propose a novel self-supervised approach, LLM2Vec-Gen, which adopts a different paradigm: rather than encoding the input, we learn to represent the model’s potential response. Specifically, we add trainable special tokens to the LLM’s vocabulary, append them to input, and optimize them to represent the LLM’s response in a fixed-length sequence. Training is guided by the LLM’s own completion for the query, along with an unsupervised embedding teacher that provides distillation targets. This formulation helps to bridge the input-output gap and transfers LLM capabilities such as safety alignment and reasoning to embedding tasks. Crucially, the LLM backbone remains frozen and training requires only unlabeled queries. LLM2Vec-Gen achieves state-of-the-art self-supervised performance on the Massive Text Embedding Benchmark (MTEB), improving by 9.3% over the best unsupervised embedding teacher. We also observe up to 43.2% reduction in harmful content retrieval and 29.3% improvement in reasoning capabilities for embedding tasks. Finally, the learned embeddings are interpretable and can be decoded into text to reveal their semantic content.
@@ 1 Introduction
Text embeddings are central to several natural language processing (NLP) applications such as semantic search, textual similarity, and retrieval-augmented generation (Reimers and Gurevych, 2019; Karpukhin et al., 2020). The success of large language models (LLMs) has led to the development of powerful LLM-based text encoders (Zhang et al., 2025b; Lee et al., 2025; BehnamGhader et al., 2024; Muennighoff et al., 2024), which leverage the web-scale pretraining of LLMs to achieve state-of-the-art results across a wide array of text encoding tasks (MTEB; Muennighoff et al., 2023; Enevoldsen et al., 2025).
@@
Although LLMs have advanced the state-of-the-art for text embedders, they still operate under the same input-centric representation paradigm adopted by BERT-based encoders (Reimers and Gurevych, 2019; Gao et al., 2021), i.e., models are trained to produce representations that capture the semantic content of the input text.111This is typically realized by taking the mean over input token representations produced by the embedding model. This paradigm struggles with the inherent nature of embedding tasks where diverse inputs are mapped to similar outputs in a shared embedding space. For example, in clustering, news articles about the same event written from different perspectives should be grouped together. Consequently, training an encoder typically requires contrastive learning on large-scale curated paired data to bridge this gap (Wang et al., 2024; Lee et al., 2025; Zhang et al., 2025b).
@@
In this work, we adopt a different paradigm for training LLM-based text encoders: rather than encoding the input itself, the model should encode the LLM’s potential response to that input (see Figure 1). This paradigm enables bridging the input-output gap and transfers LLM capabilities such as safety and reasoning into the embedding space. For instance, given a harmful query, traditional input-centric encoders represent the malicious intent of the query itself, whereas approaches adopting this paradigm encode the model’s safe refusal (e.g., “I can’t assist with that”). Similarly, reasoning capabilities which manifest in the LLM’s response rather than the input, can be transferred to the embedding space under this paradigm.
@@
We introduce LLM2Vec-Gen, as an instantiation of this paradigm. LLM2Vec-Gen is a self-supervised framework for training LLM-based text encoders to distill a model’s potential response into a fixed set of latent suffix embeddings. More specifically, given a set of unlabeled queries, we first generate responses using the LLM itself (see Figure 2). Next, we add special trainable tokens to the vocabulary and append them to the query as placeholders for the response. The last layer hidden states of these tokens are then passed through shallow projection layers to shape the response embedding. We train LLM2Vec-Gen with two complementary objectives: (1) response reconstruction, where conditioned on the embeddings, the frozen LLM reconstructs its own response in a separate forward pass, and (2) embedding alignment, where we minimize distance between the generated embedding and an unsupervised teacher’s embedding of the LLM’s response. Crucially, the LLM itself remains frozen and only the special tokens and lightweight projection layers are trained.
@@
We apply LLM2Vec-Gen to differently sized models from Llama-3.x, Qwen-2.5, and Qwen-3 model families, using corresponding unsupervised LLM2Vec (BehnamGhader et al., 2024) models as embedding teachers. We show that our proposed framework substantially outperforms existing unsupervised and self-supervised approaches on MTEB, closing over 60% of the gap to supervised methods. Notably, training with LLM2Vec-Gen consistently surpasses the off-the-shelf embedding teacher used during training, by as much as 9% for some models, which validates the effectiveness of encoding potential LLM responses. On AdvBench-IR (BehnamGhader et al., 2025), which measures retriever safety when presented with adversarial queries, models trained with LLM2Vec-Gen demonstrate superior safety, improving by 9.8 points over LLM2Vec-Qwen-3-8B. On BRIGHT (Su et al., 2025), a reasoning-intensive retrieval benchmark, LLM2Vec-Gen achieves up to 29.3% improvement over input-centric baselines, demonstrating that reasoning capabilities transfer to the embedding space. Finally, we show that our special token embeddings are interpretable and can be decoded back into text, revealing the semantic content they capture.222Our code, data, and pre-trained models are publicly available at https://github.com/McGill-NLP/llm2vec-gen
@@ 2 Related work
Recent work on embedding models has increasingly focused on repurposing decoder-only LLMs for text embedding tasks, leveraging their massive web-scale pre-training to improve performance on embedding benchmarks (Wang et al., 2024; Lee et al., 2025). Approaches such as GritLM (Muennighoff et al., 2024), which unify generation and representation within a single model, achieve strong performance on the MTEB benchmark (Muennighoff et al., 2023), but still rely on supervised contrastive learning (Khosla et al., 2020) and large-scale curated labeled datasets to align representations.
@@
To mitigate the need for labeled data, several works have explored unsupervised embedding approaches (Zhang et al., 2025a; Lin et al., 2025; Thirukovalluru and Dhingra, 2025; Jiang et al., 2024). LLM2Vec (BehnamGhader et al., 2024) shows that simple modifications like bidirectional attention and masked next-token prediction, combined with unsupervised SimCSE (Gao et al., 2021), can transform decoder-only LLMs into strong encoders. Similarly, Echo Embeddings (Springer et al., 2025), provides strong unsupervised performance, leveraging a zero-shot repetition strategy, where a model processes an “echoed” version of the input to better aggregate contextual information.
While input-centric approaches improve embedding quality, they often struggle with the lexical and conceptual gap between queries and documents. HyDE (Gao et al., 2023), one of the earliest approach under the “generate-then-encode” paradigm, demonstrates that encoding a hypothetical LLM-generated answer can yield superior retrieval performance compared to encoding the query itself. However, it incurs significant inference latency as it requires generating (long) documents during inference.
@@
More recently, several approaches focused on internalizing this generative foresight into the embedding space itself. InBedder (Peng et al., 2024) treats the instruction as a question about the input and derives embeddings from the first hidden state of the model’s generation, demonstrating that generation-derived representations can outperform prompt-based ones. GIRCSE (Tsai et al., 2026) generates soft tokens autoregressively and refines them with stepwise contrastive loss, achieving strong performance through iterative semantic refinement. Both methods rely on supervised data: InBedder requires abstractive QA pairs, while GIRCSE uses contrastive data with hard negatives. CLaRa (He et al., 2025) compresses documents into learnable memory tokens for RAG efficiency. It jointly optimizes retrieval and generation via next-token prediction loss. LLM2Vec-Gen goes beyond token-level alignment using embedding distillation to ensure embeddings semantically represent potential responses while preserving the efficiency of single-vector retrieval.
@@
Joint Embedding Predictive Architectures (JEPAs) advocate learning by predicting in representation space rather than reconstructing raw inputs (Sobal et al., 2022). Recent work extends this philosophy to language: LLM-JEPA introduces JEPA-style objectives for LLM pretraining and fine-tuning that replace or augment next-token prediction with embedding-space prediction (Huang et al., 2025). Adopting this perspective, our alignment objective shares similarities with the motivation behind JEPA: instead of matching token-level outputs, LLM2Vec-Gen predicts a target representation corresponding to the model’s likely response (via distillation from a teacher encoder), while our reconstruction objective additionally keeps the learned representations grounded in natural language and decodable.
@@ 3 LLM2Vec-Gen
The goal of LLM2Vec-Gen is to produce embeddings which represent the output the LLM would have generated for a given query, without actually generating the response at inference time. Our approach consists of the following steps (illustrated in Figure 2):
Formally, let be a pre-trained LLM and be a large corpus of queries. For every , we generate a response using . We introduce two types of new special tokens to ’s vocabulary: thought tokens () and compression tokens (). The thought tokens act as an intermediate computational buffer, while the compression tokens’ () role is to capture the semantic content of the response. For a given query , we append the thought and compression tokens and pass the combined sequence through the LLM to obtain the last layer hidden representations of the compression tokens (): , which are subsequently used for computing two objectives:
@@
The first objective ensures that the compression tokens () retain sufficient information to reconstruct the target response , i.e., we treat as a compressed representation that must preserve the content of . Concretely, we process them with a lightweight projection layer, , and feed the resulting representations as a set of soft prompts to the LLM for a second forward pass. Conditioned on this soft prompt, we train the model to reconstruct the target response via standard next-token prediction:
Embedding alignment objective.
Optimizing only the reconstruction objective might result in the soft prompts encoding the query rather than a compressed version of the target response, the goal of LLM2Vec-Gen. Consequently, we introduce a second objective to overcome this issue. For the embedding alignment objective, we utilize the encoder teacher model , which provides a target embedding . The projected representations are processed by another projection layer followed by mean pooling, , to minimize the following alignment loss:
We note that this objective is conceptually related to JEPA (Sobal et al., 2022), in that it predicts a semantic target in embedding space. Unlike JEPA, our target embedding comes from an external teacher encoder, whereas JEPA typically predicts representations produced by a learned internal target encoder.
Training and inference.
The final loss combines both objectives: . Throughout training, only the special tokens and the two MLPs are updated, while the LLM remains frozen throughout. Notably, at inference time, LLM2Vec-Gen requires only a single forward pass: we append the special tokens to the input, extract the hidden states of the compression tokens, and apply and to obtain the embedding .
@@ 4 Experiments
We evaluate LLM2Vec-Gen along three axes: general text embedding, malicious retrieval, and reasoning-intensive retrieval. After describing our setup (Sections 4.1, 4.2 and 4.3), we present results showing that LLM2Vec-Gen achieves state-of-the-art self-supervised performance and transfers LLM capabilities such as safety alignment and reasoning into the embedding space (Sections 4.4 and 4.5).
@@ 4.1 Experimental setup
Model families.
We apply LLM2Vec-Gen to decoder-only LLMs from three model families: Qwen-3 (0.6B, 1.7B, 4B, 8B; Yang et al., 2025), Qwen-2.5 (0.5B, 1.5B, 3B, 7B; Qwen et al., 2025), Llama-3.2 (1B, 3B; Meta, 2024b), and Llama-3.1 (8B; Meta, 2024a). For all models, we use 10 thought tokens () and 10 compression tokens () unless mentioned otherwise.
Encoder teacher.
This choice is guided by two requirements: 1) The teacher should share the same underlying LLM, ensuring compatible representation spaces. 2) It should have been trained with an unsupervised objective, encouraging faithful representation of the input, as opposed to contrastive objectives, which may collapse semantically distinct texts relevant to the same query (Fu et al., 2022). Hence, we use unsupervised LLM2Vec (BehnamGhader et al., 2024) models as embedding teachers.
Training.
LLM2Vec-Gen requires only an unlabeled corpus of user queries. We use single-turn questions from the Tulu instruction-following dataset (Lambert et al., 2025), sampling 160K queries from its training set. We note that the ground-truth responses from the Tulu dataset are not used for training, instead the model is trained using its own generations. We provide sample responses from different models alongside the Tulu answers in Appendix B. We train for one epoch with a batch size of 32. Training an 8B model takes approximately 3.5 hours on 2 H100 GPUs. Additional training details, including hyperparameters and compute requirements, are provided in Appendix A.
@@ 4.2 Evaluation
We evaluate LLM2Vec-Gen performance along three axes: (1) for general text embeddings, (2) for malicious retrieval, and (3) for reasoning-intensive retrieval.
General text embedding.
We evaluate LLM2Vec-Gen on MTEB(eng, v2) (Enevoldsen et al., 2025), a refined version of the Massive Text Embedding Benchmark (Muennighoff et al., 2023). MTEB(eng, v2) contains 41 tasks from seven categories: bitext mining, classification, clustering, pair classification, reranking, retrieval, and semantic textual similarity (STS). Additionally, we construct MTEB-Lite, a subset of 10 tasks that preserves the category distribution of the full benchmark (see Appendix A). We use MTEB-Lite for ablations and analysis. We report the average score across all tasks as the primary metric.
Malicious retrieval.
We additionally evaluate on AdvBench-IR (BehnamGhader et al., 2025), a benchmark that measures retriever vulnerability to malicious queries. The benchmark contains 520 harmful queries derived from AdvBench (Zou et al., 2023) spanning five harm categories: cybercrime, chemical and biological weapons, misinformation, harassment, and illegal activities. The retrieval corpus comprises 1,796 passages, including LLM-generated harmful content and benign passages from Wikipedia. We report top-5 accuracy, which measures how often a harmful passage appears among the top 5 retrieved results. Lower scores indicate safer retrieval behavior.
Reasoning-intensive retrieval.
Lastly, we evaluate on BRIGHT (Su et al., 2025), which is specifically designed to assess retrieval capabilities in scenarios requiring intensive reasoning to determine the relevance between query and document. Instead of surface-level semantic matching, BRIGHT consists of real-world queries across diverse domains including biology, coding, math, and physics, where the relationship between the query and the relevant document requires logical deduction. We focus on the zero-shot retrieval performance, reporting nDCG@10.
Text embedders typically prepend inputs with task-specific instructions, such as ‘Retrieve text that answers this query’ or ‘Retrieve text that is semantically similar to this text’. We reformulate these as generative instructions, e.g., ‘Generate text that answers this query’ and ‘Generate text that is semantically similar to this text’ and employ a generic summarization instruction for documents: ‘Summarize the following passage’. Refer to Tables 7 and 8 for the instructions used in all our evaluation datasets.

### v2

@@ Abstract
Fine-tuning LLM-based text embedders via contrastive learning maps inputs and outputs into a new representational space, discarding the LLM’s output semantics. We propose LLM2Vec-Gen, a self-supervised alternative that instead produces embeddings directly in the LLM’s output space by learning to represent the model’s potential response. Specifically, trainable special tokens are appended to the input and optimized to compress the LLM’s own response into a fixed-length embedding, guided by an unsupervised embedding teacher and a reconstruction objective. Crucially, the LLM backbone remains frozen and training requires only unlabeled queries. LLM2Vec-Gen achieves state-of-the-art self-supervised performance on the Massive Text Embedding Benchmark (MTEB), improving by 8.8% over the unsupervised embedding teacher. Since the embeddings preserve the LLM’s response-space semantics, they inherit capabilities such as safety alignment (up to 22.6% reduction in harmful content retrieval) and reasoning (up to 35.6% improvement on reasoning-intensive retrieval). Finally, the learned embeddings are also interpretable: they can be decoded back into text to reveal their semantic content.
@@ 1 Introduction
Text embeddings are central to several natural language processing tasks such as semantic search, textual similarity, and retrieval-augmented generation (Reimers and Gurevych, 2019; Karpukhin et al., 2020). The success of large language models (LLMs) has led to powerful LLM-based text encoders (Zhang et al., 2025b; Lee et al., 2025; BehnamGhader et al., 2024; Muennighoff et al., 2024), which leverage pretrained LLMs to achieve state-of-the-art results across embedding benchmarks (MTEB; Muennighoff et al., 2023; Enevoldsen et al., 2025).
@@
Yet these models still follow the input-centric paradigm of BERT-based encoders (Reimers and Gurevych, 2019; Gao et al., 2021): they represent the semantic content of the input text, typically by taking the mean over input token representations produced by the embedding model. Because embedding tasks often require mapping diverse inputs to similar outputs (e.g., clustering articles about the same event), this paradigm relies on contrastive learning over curated paired data to project both queries and documents into a new shared embedding space (Wang et al., 2024; Lee et al., 2025; Zhang et al., 2025b). Crucially, this new space no longer coincides with the model’s response representations, discarding the rich semantic structure the LLM acquired during pretraining.
@@
In this work, we adopt a fundamentally different paradigm: rather than encoding the input, the model should encode the LLM’s potential response to that input (see Figure 1). By keeping embeddings closer to the LLM’s response space, this paradigm preserves capabilities that manifest in the model’s responses. For instance, given a harmful query, input-centric encoders represent the malicious intent, whereas our approach encodes the model’s safe refusal (e.g., “I cannot assist with that”). Similarly, reasoning capabilities that emerge in the LLM’s responses transfer naturally to the embedding space.
@@
We introduce LLM2Vec-Gen, a self-supervised instantiation of this paradigm that distills the model’s potential response into a fixed set of latent suffix embeddings (see Figure 2). Given a set of unlabeled queries, we first generate responses using the LLM itself. Next, we add special trainable tokens to the vocabulary and append them to the query as placeholders for the response. The hidden states of these tokens are passed through lightweight projection layers to produce the response embedding, trained with two complementary objectives: (1) embedding alignment, where input’s embedding is aligned with an unsupervised teacher’s embedding of the LLM’s response, and (2) response reconstruction, where the frozen LLM reconstructs its own response conditioned on the embeddings. The LLM backbone remains frozen and only the special tokens and projection layers are trained. Moreover, LLM2Vec-Gen is entirely self-supervised: it requires only unlabeled queries as training data, utilizing the backbone LLM itself for response generation and the unsupervised embedding teacher, eliminating the need for curated paired data across different stages of the training.
@@
We apply LLM2Vec-Gen to models from Llama-3.x, Qwen-2.5, and Qwen-3 families, using the corresponding unsupervised LLM2Vec (BehnamGhader et al., 2024) models as embedding teachers. LLM2Vec-Gen substantially outperforms existing unsupervised and self-supervised methods on MTEB, closing over 60% of the gap to supervised methods and consistently surpassing the embedding teacher by up to 8.8%, validating the effectiveness of output-centric representations. On AdvBench-IR (BehnamGhader et al., 2025), which measures retriever safety when presented with adversarial queries, LLM2Vec-Gen improves safety by 9.2 points over the embedding teacher. On the reasoning-intensive BRIGHT benchmark (Su et al., 2025), LLM2Vec-Gen achieves up to 35.6% improvement over the input-centric baseline, demonstrating that reasoning capabilities transfer to the embedding space. Finally, we show that the learned embeddings are interpretable and can be decoded back into text, revealing the semantic content they capture.111Our code, data, and pretrained models are publicly available at https://github.com/McGill-NLP/llm2vec-gen.
@@ 2 Related work
Recent work has repurposed decoder-only LLMs for text embedding, leveraging their massive web-scale pretraining to improve performance on embedding benchmarks (Lee et al., 2024; Wang et al., 2024; Lee et al., 2025). Approaches such as GritLM (Muennighoff et al., 2024) unify generation and representation but still rely on supervised contrastive learning (Khosla et al., 2020) and large-scale curated labeled datasets to align representations.
@@
To mitigate the need for labeled data, several works have explored unsupervised embedding approaches (Zhang et al., 2025a; Lin et al., 2025; Thirukovalluru and Dhingra, 2025; Jiang et al., 2024; Lei et al., 2024). LLM2Vec (BehnamGhader et al., 2024) shows that simple modifications like bidirectional attention and masked next-token prediction, combined with unsupervised SimCSE (Gao et al., 2021), can transform decoder-only LLMs into strong encoders. Similarly, Echo Embeddings (Springer et al., 2025) uses an input repetition strategy for effective embedding. All these unsupervised methods remain input-centric: they represent what the text says rather than what the LLM would respond, and they often struggle with the lexical and conceptual gap between queries and documents.
Several recent approaches employ learnable latent or compression tokens to produce compact representations within LLMs. xRAG (Cheng et al., 2024) compresses retrieved document into a single latent token and projects it into the language model’s representation space for efficient RAG. CLaRa (He et al., 2025) compresses documents into learnable memory tokens and jointly optimizes retrieval and generation end-to-end via next-token prediction loss. While these methods leverage special tokens for compression, they remain fundamentally input-centric — compressing or summarizing the given text. In contrast, LLM2Vec-Gen uses trainable compression tokens to model the LLM’s potential response.
@@
A few recent works have explored output-centric embeddings. HyDE (Gao et al., 2023) demonstrated the value of encoding LLM-generated answers rather than queries, but requires generating multiple answers at inference time. More recent methods internalize this generative foresight into the embedding space itself: InBedder (Peng et al., 2024) derives embeddings from the first generated hidden state using abstractive QA supervision, demonstrating that generation-derived representations can outperform prompt-based ones. GIRCSE (Tsai et al., 2026) generates soft tokens autoregressively and refines them with stepwise contrastive loss using hard negatives. Both methods rely on supervised data: InBedder requires abstractive QA pairs, while GIRCSE uses contrastive data with hard negatives. Unlike these approaches, LLM2Vec-Gen uses embedding-level distillation from an unsupervised teacher to ensure embeddings faithfully represent the LLM’s potential response, without requiring supervised data or inference-time autoregressive generation.
@@
Our alignment objective also connects to Joint Embedding Predictive Architectures (JEPAs; Sobal et al., 2022), which advocate predicting in representation space rather than reconstructing raw inputs, recently extended to language by LLM-JEPA (Huang et al., 2025). Adopting this perspective, LLM2Vec-Gen predicts a target representation of the model’s likely response via external teacher distillation, while the reconstruction objective keeps the learned representations grounded in natural language.
@@ 3 LLM2Vec-Gen
The goal of LLM2Vec-Gen is to produce embeddings that represent the output the LLM would have generated for a given query, without actually generating the response at inference time. Our training recipe consists of the following steps (illustrated in Figure 2):
Given a dataset of queries, we generate target responses from the LLM.
We use an off-the-shelf teacher embedding model to construct embeddings for the generated responses.
Special tokens are added to the model’s vocabulary and appended to input queries.
Two losses are computed by (a) comparing the special tokens’ final representations to the teacher’s response embedding, and (b) conditioning the LLM on the special tokens and computing a cross-entropy loss with the response as a language modeling target.
Formally, let be a pretrained LLM and be a large corpus of queries. For every , we generate a response using . We introduce new special tokens to ’s vocabulary: . The role of these compression tokens is to capture the semantic content of the response. For a given query , we append the compression tokens and pass the combined sequence through the LLM to obtain the last layer hidden representations of the compression tokens only: , which are subsequently used for computing embedding alignment and reconstruction objectives.
Embedding alignment objective.
LLM2Vec-Gen’s primary goal is to embed queries into the LLM’s response space. Consequently, we introduce the embedding alignment objective, where we utilize an unsupervised encoder teacher model , to provide a target embedding for a generated response . The hidden representations of compression tokens are projected through two lightweight projection layers222The second projection layer is required for distillation from embedding teacher models with different hidden dimensions. followed by mean pooling, , to minimize the following mean squared loss (Reimers and Gurevych, 2020):
A key distinction of LLM2Vec-Gen from standard contrastive learning is that contrastive objectives map queries and documents into a new shared latent space using relative relevance judgments, often with hard negatives. In contrast, our teacher uses only SimCSE (Gao et al., 2021), which enforces uniformity by pushing apart random negatives, largely preserving the LLM’s local representational geometry — unlike supervised contrastive learning, which restructures the space around relevance labels. By distilling from positive LLM-generated responses rather than discriminative pairs, LLM2Vec-Gen encourages embeddings to reflect the semantic content of the LLM’s responses, inheriting properties such as safety and reasoning (Section 4.5).
@@
Our goal is to generate embeddings that not only encode the potential response of the underlying LLM but are also interpretable, i.e., they can be decoded by the LLM to reveal their content. To this end, the second objective ensures that the compression tokens () retain sufficient information to reconstruct the target response . Concretely, we project the compression tokens’ hidden representations () using a projection layer, and feed the resulting representations () as a set of soft prompts to the LLM for a second forward pass. Conditioned on this soft prompt, we train the model to reconstruct the target response via standard next-token prediction:
This objective forces to serve as an information bottleneck, compressing the content of (Cheng et al., 2024; He et al., 2025).
Training and inference. The final loss combines both objectives: . Throughout training, only the special tokens and the two MLPs are updated, while the LLM remains frozen. Notably, at inference time, LLM2Vec-Gen requires only a single forward pass: we append the special tokens to the input, extract the hidden states of the compression tokens, and apply two MLPs to obtain the embedding . The second forward pass can be an optional extra step to reveal the content of the embeddings.
@@ 4 Experiments
We evaluate LLM2Vec-Gen along three axes: general text embedding, malicious retrieval, and reasoning-intensive retrieval. After describing our setup (Sections 4.1, 4.2 and 4.3), we show that LLM2Vec-Gen achieves state-of-the-art self-supervised performance and transfers LLM capabilities such as safety and reasoning into the embedding space (Sections 4.4 and 4.5).
@@ 4.1 Experimental setup
Model families. We apply LLM2Vec-Gen to decoder-only LLMs from three model families: Qwen-3 (0.6B, 1.7B, 4B, 8B; Yang et al., 2025), Qwen-2.5 (0.5B, 1.5B, 3B, 7B; Qwen et al., 2025), Llama-3.2 (1B, 3B; Meta, 2024b), and Llama-3.1 (8B; Meta, 2024a). For all models, we use 10 compression tokens () unless mentioned otherwise.
Encoder teacher. This choice is guided by two requirements: (1) The teacher should share the same underlying LLM, ensuring compatible representation spaces. (2) It should be trained without labeled data, so that it produces faithful content representations rather than relevance-biased ones (Fu et al., 2022). As discussed in Section 3, the teacher’s unsupervised objective applies only light uniformity regularization, preserving the LLM’s representational geometry and ensuring the alignment target faithfully represents its input response content. We use unsupervised LLM2Vec (BehnamGhader et al., 2024) models as embedding teachers. These criteria ensure that LLM2Vec-Gen remains fully self-supervised.
Training. LLM2Vec-Gen requires only an unlabeled corpus of user queries. We use 160K single-turn questions from the Tulu instruction-following dataset (Lambert et al., 2025): ground-truth responses are not used — the model is trained on its own generations (sample responses in Appendix D). We train for one epoch with a batch size of 32; an 8B model takes approximately 3.5 hours on 2 H100 GPUs. See additional training details in Appendix C.
@@ 4.2 Evaluation
We evaluate LLM2Vec-Gen performance along three axes: (1) general text embeddings, (2) malicious retrieval, and (3) reasoning-intensive retrieval.
General text embedding. We evaluate LLM2Vec-Gen on MTEB(eng, v2) (Enevoldsen et al., 2025), which contains 41 tasks across seven categories: bitext mining, classification, clustering, pair classification, reranking, retrieval, and semantic textual similarity (STS). We report the average score across all tasks. Additionally, we construct MTEB-Lite, a subset of 10 tasks that preserves the category distribution of the full benchmark (see Appendix C). We use MTEB-Lite only for ablations.
Malicious retrieval. We additionally evaluate on AdvBench-IR (BehnamGhader et al., 2025), a benchmark that measures retriever vulnerability to malicious queries. The benchmark contains 520 harmful queries derived from AdvBench (Zou et al., 2023) spanning five harm categories: cybercrime, chemical and biological weapons, misinformation, harassment, and illegal activities. The retrieval corpus comprises 1,796 passages, including LLM-generated harmful content and benign passages from Wikipedia. We report top-5 accuracy; lower scores indicate safer retrieval behavior.
Reasoning-intensive retrieval. Lastly, we evaluate on BRIGHT (Su et al., 2025), which assesses retrieval in scenarios requiring intensive reasoning to determine query-document relevance. Instead of surface-level semantic matching, BRIGHT consists of real-world queries across diverse domains, including biology, coding, math, and physics, where relevance requires logical deduction. We report nDCG@10 for zero-shot retrieval performance.
Since LLM2Vec-Gen encodes outputs rather than inputs, we reformulate standard task instructions from embedding-oriented (e.g., ‘Retrieve text that answers this query’ or ‘Retrieve text that is semantically similar to this text’) to generative (e.g., ‘Generate text that answers this query’ and ‘Generate text that is semantically similar to this text’) and use a summarization instruction for documents: ‘Summarize the following passage’. Refer to Tables 7 and 8 for the instructions used in all our evaluation datasets. Figure 5 validates this design choice.